OpenStudy (anonymous):

True or Falus The Inverse of f(x) = 2x - 10 is f^ -1 (x) = 1/2 x +5

OpenStudy (ivettef365):

y = 2x - 10 y + 10 = 2x x = y + 10 ------ 2 x = 1/2y + 5

OpenStudy (ivettef365):

now switch the y for x

OpenStudy (ivettef365):

so the inverse = 1/2x + 5

OpenStudy (loser66):

you can check by f(f^-1)=x. therefore, f^-1 is inverse. I mean f(f^1)= 2(1/2x +5) -10 =x therefore,1/2 x +5 is inverse of 2x -10

OpenStudy (anonymous):

Its True Thanks for your help everyone